# Minimal model program for log canonical pairs on complex analytic spaces

### v1

@@ 3.2. Definition of log MMP
    • $\phi\colon X\dashrightarrow X^{\prime}$ is a bimeromorphic contraction and $f\colon X\to Z$ and $f^{\prime}\colon X^{\prime}\to Z$ are bimeromorphisms,
    • $f$ is a contraction of a $(K_{X}+\Delta)$-negative extremal ray of $\overline{\rm NE}(X/Y;W)$, in particular, $\rho(X/Y;W)-\rho(Z/Y;W)=1$ (cf. [F22b, Theorem 1.2 (4)(iii)]) and $-(K_{X}+\Delta)$ is ample over $Z$, and
@@
    For the simplicity of notation, a sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ is denoted by
    $(X_{0},\Delta_{0})\dashrightarrow(X_{1},\Delta_{1})\dashrightarrow\cdots\dashrightarrow(X_{i},\Delta_{i})\dashrightarrow\cdots.$
@@
    Let $H$ be an $\mathbb{R}$-Cartier divisor on $X$. We put
    $H_{0}:=H|_{X_{0}}\qquad{\rm and}\qquad H_{i+1}:=(\phi_{i*}H_{i})|_{X_{i+1}}$
    for each $i\geq 0$. Then a sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ with scaling of $H$ is a sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$
    with the data $\{Y_{i}\}_{i\geq 0}$ and $\{\phi_{i}\colon X_{i}\dashrightarrow X^{\prime}_{i}\}_{i\geq 0}$ such that $(X_{i},\Delta_{i})\dashrightarrow(X^{\prime}_{i},\phi_{i*}\Delta_{i})$ is a step of a $(K_{X_{i}}+\Delta_{i})$-MMP over $Y$ around $W$ with scaling of $H_{i}$ for every $i$.
@@
be a sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ defined with $\{Y_{i}\}_{i\geq 0}$ and $\{\phi_{i}\colon X_{i}\dashrightarrow X^{\prime}_{i}\}_{i\geq 0}$. We say that the $(K_{X}+\Delta)$-MMP is represented by bimeromorphic contractions if $Y_{i}=Y$ for all $i\geq 0$.
@@ 3.3. Basics of log MMP
[…]
@@
By the ampleness of $-(K_{X}+\Delta)$ over $Z$ and [F22b, Theorem 9.1], after shrinking $Y$ around $W$, we can find $m_{0}\in\mathbb{Z}_{>0}$ such that $mD$ is base point free for all $m\geq m_{0}$. Let $X\to Z_{1}$ and $X\to Z_{2}$ be the contractions over $Z$ induced by $(m_{0}+1)D$ and $(m_{0}+2)D$, respectively. Since $D\equiv_{W}0$, the induced morphisms $Z_{1}\to Z$ and $Z_{2}\to Z$ are biholomorphisms near the inverse image of $W$. By shrinking $Y$ around $W$ if necessary, we may assume $Z_{1}=Z_{2}=Z$. Then there are Cartier divisors $H_{1}$ and $H_{2}$ on $Z$ such that $(m_{0}+1)D\sim f^{*}H_{1}$ and $(m_{0}+2)D\sim f^{*}H_{2}$. Then
@@
[…]
@@
Since $\rho(X/Y;W)-\rho(Z/Y;W)=1$ and $-(K_{X}+\Delta)$ is ample over $Z$, by shrinking $Y$ around $W$, we can find $r\in\mathbb{R}$ and an $\mathbb{R}$-Cartier divisor $L$ on $Z$ such that
@@ 3.4. Nef thresholds in log MMP
[…]
@@
such that $D_{j}$ are Cartier and trivially intersects any curve over $W$ contracted by $f_{i}$. By the ampleness of $-(K_{X_{i}}+\Delta_{i})$ over $Z_{i}$ and [F22b, Theorem 9.1], after shrinking $Y$ around $W$, it follows that $D_{j}$ are semi-ample over $Z_{i}$ for all $j$. Let $X_{i}\to V_{j}$ be the contractions over $Z_{i}$ induced by $D_{j}$. By the above condition for $D_{j}$, the induced morphisms $V_{j}\to Z_{i}$ are biholomorphisms near the inverse image of $W$. By shrinking $Y$ around $W$, we may assume $V_{j}=Z_{i}$ for all $j$. Then there are $\mathbb{Q}$-Cartier divisors $H_{j}$ on $Z_{i}$ such that $D_{j}\sim_{\mathbb{Q}}f_{i}^{*}H_{j}$ for all $j$. Put $A:=\sum_{j}r_{j}H_{j}$. Then
@@ 3.5. From existence of weak lc model to termination
(***) $\begin{split}K_{\widetilde{X}}+\widetilde{\Delta}+t\widetilde{H}=f^{*}(K_{X}+\Delta+tH)+\widetilde{E}_{t}.\end{split}$
@@
(***** *) $\begin{split}K_{\widetilde{X}}+\widetilde{\Delta}=f^{\prime*}(K_{X^{\prime}}+\Delta^{\prime})+\widetilde{E}_{0}+\widetilde{F}.\end{split}$
@@
${\rm coeff}_{P}(\widetilde{E}_{0})=a(P,X,\Delta)+{\rm coeff}_{P}(\widetilde{\Delta}).$
@@
$-1=a(P,X^{\prime},\Delta^{\prime})\geq a(P,X,\Delta)\geq-1,$
@@
In this step, we run a $(K_{\widetilde{X}}+\widetilde{\Delta})$-MMP by using Lemma 3.14 and construct an lc pair $(\widetilde{X}^{\prime},\widetilde{\Delta}^{\prime})$ with a projective bimeromorphism $\widetilde{X}^{\prime}\to X^{\prime}$.
Let $\pi^{\prime}\colon X^{\prime}\to Y$ be the structure morphism. Let $L^{\prime}$ be a $\pi^{\prime}$-ample Cartier divisor on $X^{\prime}$, and let $\widetilde{L}\geq 0$ be an $\mathbb{R}$-divisor on $\widetilde{X}$ such that $(\widetilde{X},\widetilde{\Delta}+\widetilde{H}+\widetilde{L})$ is an lc pair and $\widetilde{L}\sim_{\mathbb{R}}(3\cdot{\rm dim}\,X)f^{\prime*}L^{\prime}$. We run a $(K_{\widetilde{X}}+\widetilde{\Delta}+\widetilde{L})$-MMP over $Y$ around $W$ with scaling of a $(\pi^{\prime}\circ f^{\prime})$-ample divisor
@@
Since $(X,\Delta)$ has a weak lc model over $Y$ around $W$, $K_{X}+\Delta$ is $\pi$-pseudo-effective. Thus $K_{\widetilde{X}}+\widetilde{\Delta}+\widetilde{L}$ is $(\pi^{\prime}\circ f^{\prime})$-pseudo-effective. By [F22a, Lemma 13.7], there is $j^{\prime}\gg 0$ such that $K_{\widetilde{X}_{j^{\prime}}}+\widetilde{\Delta}_{j^{\prime}}+\widetilde{L}_{j^{\prime}}\in\overline{\rm Mov}(\widetilde{X}_{j^{\prime}}/Y;W)$. After shrinking $Y$ around $W$, we may assume that the $(K_{\widetilde{X}}+\widetilde{\Delta}+\widetilde{L})$-MMP $(\widetilde{X}_{0},\widetilde{\Delta}_{0}+\widetilde{L}_{0})\dashrightarrow(\widetilde{X}_{j^{\prime}},\widetilde{\Delta}_{j^{\prime}}+\widetilde{L}_{j^{\prime}})$ is a bimeromorphic contraction. Put
@@
By Lemma 3.14, the induced bimeromorphic map $\widetilde{f}^{\prime}\colon\widetilde{X}^{\prime}\dashrightarrow X^{\prime}$ is a morphism. Let $\widetilde{E}^{\prime}$ and $\widetilde{F}^{\prime}$ be the strict transforms of $\widetilde{E}_{0}$ and $\widetilde{F}$ on $\widetilde{X}^{\prime}$, respectively. Then we have $\widetilde{L}^{\prime}\sim_{\mathbb{R}}(3\cdot{\rm dim}\,X)\widetilde{f}^{\prime*}L^{\prime}$, and (***** *1) implies
@@
Therefore $K_{\widetilde{X}^{\prime}}+\widetilde{\Delta}^{\prime}\in\overline{\rm Mov}(\widetilde{X}^{\prime}/X^{\prime};\pi^{\prime-1}(W))$. Moreover, $\widetilde{E}^{\prime}+\widetilde{F}^{\prime}$ is $\widetilde{f}^{\prime}$-exceptional by construction. By the negativity lemma (Corollary 2.15), after shrinking $Y$ around $W$ we have $\widetilde{E}^{\prime}+\widetilde{F}^{\prime}=0$. Therefore, $K_{\widetilde{X}^{\prime}}+\widetilde{\Delta}^{\prime}=\widetilde{f}^{\prime*}(K_{X^{\prime}}+\Delta^{\prime})$ is nef over $W$. Now recall by Lemma 3.14 that the $(K_{\widetilde{X}}+\widetilde{\Delta}+\widetilde{L})$-MMP over $Y$ around $W$ is also a $(K_{\widetilde{X}}+\widetilde{\Delta})$-MMP over $Y$ around $W$. Hence, $(\widetilde{X}^{\prime},\widetilde{\Delta}^{\prime})$ is a log minimal model of $(\widetilde{X},\widetilde{\Delta})$ over $Y$ around $W$.
@@
Let $\widetilde{H}^{\prime}$ be the strict transform of $\widetilde{H}$ on $\widetilde{X}^{\prime}$. We take a positive real number $\epsilon$ such that $\widetilde{X}\dashrightarrow\widetilde{X}^{\prime}$ is a sequence of steps of a $(K_{\widetilde{X}}+\widetilde{\Delta}+\epsilon\widetilde{H})$-MMP over $Y$ around $W$. By Lemma 3.10 and taking $\epsilon$ sufficiently small, we may assume that for any sequence of steps of a $(K_{\widetilde{X}^{\prime}}+\widetilde{\Delta}^{\prime}+\epsilon\widetilde{H}^{\prime})$-MMP over $Y$ around $W$
@@
$K_{\widetilde{X}^{\prime}_{k}}+\widetilde{\Delta}^{\prime}_{k}$ trivially intersects the curves over $W$ that are contracted by the $(k+1)$-th steps of the MMP. In particular, for every $k\geq 0$, after shrinking $Y$ around $W$, it follows that $(\widetilde{X}^{\prime}_{k},\widetilde{\Delta}^{\prime}_{k})$ is a weak lc model of $(\widetilde{X}^{\prime},\widetilde{\Delta}^{\prime})$ over $Y$ around $W$.
@@
is represented by bimeromorphic contractions and $X=X_{0}$.
@@
${(\widetilde{X},\widetilde{\Delta})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${(\widetilde{X}^{\prime},\widetilde{\Delta}^{\prime})}$${(X,\Delta)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(X_{n},\Delta_{n})}$
@@
    $f\colon\widetilde{X}\to X$ is a log resolution of $(X,\Delta+H)$ such that for any $t\in[0,1]$, we can write
    $K_{\widetilde{X}}+\widetilde{\Delta}+t\widetilde{H}=f^{*}(K_{X}+\Delta+tH)+\widetilde{E}_{t}$
    for some effective $f$-exceptional $\mathbb{R}$-divisor $\widetilde{E}_{t}$ on $\widetilde{X}$, where $\widetilde{H}=f^{-1}_{*}H$,
@@
    $\widetilde{X}\dashrightarrow\widetilde{X}^{\prime}$ is a bimeromorphic contraction that defines:
    • a finite sequence of steps of a $(K_{\widetilde{X}}+\widetilde{\Delta})$-MMP over $Y$ around $W$
      $(\widetilde{X},\widetilde{\Delta})\dashrightarrow(\widetilde{X}^{\prime},\widetilde{\Delta}^{\prime})$
      terminating with a log minimal model $(\widetilde{X}^{\prime},\widetilde{\Delta}^{\prime})$ over $Y$ around $W$, and
    • a finite sequence of steps of a $(K_{\widetilde{X}}+\widetilde{\Delta}+\epsilon\widetilde{H})$-MMP over $Y$ around $W$
      $(\widetilde{X},\widetilde{\Delta}+\epsilon\widetilde{H})\dashrightarrow(\widetilde{X}^{\prime},\widetilde{\Delta}^{\prime}+\epsilon\widetilde{H}^{\prime}),$
@@
    for any sequence of steps of a $(K_{\widetilde{X}^{\prime}}+\widetilde{\Delta}^{\prime}+\epsilon\widetilde{H}^{\prime})$-MMP over $Y$ around $W$
    $(\widetilde{X}^{\prime}_{0},\widetilde{\Delta}^{\prime}_{0}+\epsilon\widetilde{H}^{\prime}_{0})\dashrightarrow(\widetilde{X}^{\prime}_{1},\widetilde{\Delta}^{\prime}_{1}+\epsilon\widetilde{H}^{\prime}_{1})\dashrightarrow\cdots\dashrightarrow(\widetilde{X}^{\prime}_{k},\widetilde{\Delta}^{\prime}_{k}+\epsilon\widetilde{H}^{\prime}_{k})\dashrightarrow\cdots$
    and any $k\geq 0$, after shrinking $Y$ around $W$ suitably, $(\widetilde{X}^{\prime}_{k},\widetilde{\Delta}^{\prime}_{k})$ is a weak lc model of $(\widetilde{X}^{\prime},\widetilde{\Delta}^{\prime})$ over $Y$ around $W$, and
@@
    $X\dashrightarrow X_{n}$ is a bimeromorphic contraction that defines a finite sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ with scaling of $H$
    $(X,\Delta)\dashrightarrow(X_{n},\Delta_{n})$
    such that $\lambda_{n}<{\rm min}\{\lambda_{n-1},\epsilon\}$.
@@
In this step, we will prove that after shrinking $Y$ around $W$ suitably, $(X_{n},\Delta_{n}+\lambda^{\prime}H_{n})$ is a weak lc model of $(\widetilde{X}^{\prime},\widetilde{\Delta}^{\prime}+\lambda^{\prime}\widetilde{H}^{\prime})$ over $Y$ around $W$.
By definition and (i) in Step 3, $(X_{n},\Delta_{n}+\lambda^{\prime}H_{n})$ is a weak lc model of $(\widetilde{X},\widetilde{\Delta}+\lambda^{\prime}\widetilde{H})$ over $Y$ around $W$. Moreover, since $\lambda^{\prime}\leq\epsilon$, by (ii) in Step 3, the bimeromorphic map $\widetilde{X}\dashrightarrow\widetilde{X}^{\prime}$ defines a sequence of steps of a $(K_{\widetilde{X}}+\widetilde{\Delta}+\lambda^{\prime}\widetilde{H})$-MMP over $Y$ around $W$
$(\widetilde{X},\widetilde{\Delta}+\lambda^{\prime}\widetilde{H})\dashrightarrow(\widetilde{X}^{\prime},\widetilde{\Delta}^{\prime}+\lambda^{\prime}\widetilde{H}^{\prime}).$
By Lemma 3.11, after shrinking $Y$ around $W$ suitably, $(X_{n},\Delta_{n}+\lambda^{\prime}H_{n})$ is a weak lc model of $(\widetilde{X}^{\prime},\widetilde{\Delta}^{\prime}+\lambda^{\prime}\widetilde{H}^{\prime})$ over $Y$ around $W$.
@@
Since ${\rm Supp}\,(\widetilde{\Delta}+\epsilon\widetilde{H})$ contains all $f$-exceptional divisors and components of $f^{-1}_{*}(\Delta+H)$, we may assume that $\widetilde{\Gamma}\geq 0$. By (i) in Step 3, we have
@@
Since $\Gamma$ is general, we may assume that $(\widetilde{X},\widetilde{\Gamma})$ is log smooth and lc, and $\Gamma\sim_{\mathbb{R}}\delta\Delta+\epsilon H$ shows
@@
Then $(\widetilde{X},\widetilde{\Gamma})$ is klt. Indeed, if a prime divisor $Q$ on $\widetilde{X}$ is not a component of $\widetilde{E}_{\epsilon}$, then
@@
Since $(\widetilde{X},\widetilde{\Gamma})$ is log smooth, we see that $(\widetilde{X},\widetilde{\Gamma})$ is klt.
Now $\widetilde{\Gamma}$ is $(\pi\circ f)$-big. Since $\gamma>0$ is sufficiently small, we may assume that $\widetilde{X}\dashrightarrow\widetilde{X}^{\prime}$ is a finite sequence of steps of a $(K_{\widetilde{X}}+\widetilde{\Delta}+\epsilon\widetilde{H}-\gamma\widetilde{E}_{\epsilon})$-MMP over $Y$ around $W$. Here, we used the second part of (ii) in Step 3. Then $\widetilde{X}\dashrightarrow\widetilde{X}^{\prime}$ is also a $(K_{\widetilde{X}}+\widetilde{\Gamma})$-MMP over $Y$ around $W$. Let $\widetilde{\Gamma}^{\prime}$ be the strict transform of $\widetilde{\Gamma}$ on $\widetilde{X}^{\prime}$. Then $(\widetilde{X}^{\prime},\widetilde{\Gamma}^{\prime})$ is klt. Moreover, by the argument of Step 2, the strict transform of $\widetilde{E}_{0}$ on $\widetilde{X}^{\prime}$ is zero. Since $\widetilde{E}_{\epsilon}\leq\widetilde{E}_{0}$, the strict transform of $\widetilde{E}_{\epsilon}$ on $\widetilde{X}^{\prime}$ is zero. This implies
@@
Since $\widetilde{\Gamma}^{\prime}$ is big over $Y$, by [F22a, Theorem 1.1] and shrinking $Y$ around $W$, we get a bimeromorphic contraction $\widetilde{X}^{\prime}\dashrightarrow\widetilde{X}^{\prime\prime}$ over $Y$ that defines a finite sequence of steps of a $(K_{\widetilde{X}^{\prime}}+\widetilde{\Gamma}^{\prime})$-MMP over $Y$ around $W$ terminating with a log minimal model $(\widetilde{X}^{\prime\prime},\widetilde{\Gamma}^{\prime\prime})$. Let $\widetilde{\Delta}^{\prime\prime}$ (resp. $\widetilde{H}^{\prime\prime}$) be the strict transform of $\widetilde{\Delta}^{\prime}$ (resp. $\widetilde{H}^{\prime}$) on $\widetilde{X}^{\prime\prime}$. Then $\widetilde{X}^{\prime}\dashrightarrow\widetilde{X}^{\prime\prime}$ defines a sequence of steps of a $(K_{\widetilde{X}^{\prime}}+\widetilde{\Delta}^{\prime}+\epsilon\widetilde{H}^{\prime})$-MMP over $Y$ around $W$ terminating with a log minimal model $(\widetilde{X}^{\prime\prime},\widetilde{\Delta}^{\prime\prime}+\epsilon\widetilde{H}^{\prime\prime})$ over $Y$ around $W$. Therefore, we have a bimeromorphic contraction
@@
that defines a sequence of steps of a $(K_{\widetilde{X}}+\widetilde{\Delta}+\epsilon\widetilde{H})$-MMP over $Y$ around $W$ terminating with a log minimal model $(\widetilde{X}^{\prime\prime},\widetilde{\Delta}^{\prime\prime}+\epsilon\widetilde{H}^{\prime\prime})$ over $Y$ around $W$. By (iii) in Step 3, $(\widetilde{X}^{\prime\prime},\widetilde{\Delta}^{\prime\prime})$ is a weak lc model of $(\widetilde{X},\widetilde{\Delta})$ over $Y$ around $W$. Since $0\leq\lambda_{n}<\lambda^{\prime}\leq\epsilon$ by (iv) in Step 3, we see that $(\widetilde{X}^{\prime\prime},\widetilde{\Delta}^{\prime\prime}+\lambda^{\prime}\widetilde{H}^{\prime\prime})$ and $(\widetilde{X}^{\prime\prime},\widetilde{\Delta}^{\prime\prime}+\lambda_{n}\widetilde{H}^{\prime\prime})$ are weak lc models of $(\widetilde{X},\widetilde{\Delta}+\lambda^{\prime}\widetilde{H})$ and $(\widetilde{X},\widetilde{\Delta}+\lambda_{n}\widetilde{H})$ over $Y$ around $W$, respectively.
@@
By (i) and (iv) in Step 3 and the definition of weak lc models, $(X_{n},\Delta_{n}+\lambda^{\prime}H_{n})$ and $(X_{n},\Delta_{n}+\lambda_{n}H_{n})$ are also weak lc models of $(\widetilde{X},\widetilde{\Delta}+\lambda^{\prime}\widetilde{H})$ and $(\widetilde{X},\widetilde{\Delta}+\lambda_{n}\widetilde{H})$ over $Y$ around $W$, respectively. Let $g\colon V\to X_{n}$ and $\widetilde{g}\colon V\to\widetilde{X}^{\prime\prime}$ be a common resolution of the induced bimeromorphic map $X_{n}\dashrightarrow\widetilde{X}^{\prime\prime}$. By Lemma 3.3 and shrinking $Y$ around $W$, we have
$\begin{split}g^{*}(K_{X_{n}}+\Delta_{n}+\lambda^{\prime}H_{n})&=\widetilde{g}^{*}(K_{\widetilde{X}^{\prime\prime}}+\widetilde{\Delta}^{\prime\prime}+\lambda^{\prime}\widetilde{H}^{\prime\prime}),\quad{\rm and}\\ g^{*}(K_{X_{n}}+\Delta_{n}+\lambda_{n}H_{n})&=\widetilde{g}^{*}(K_{\widetilde{X}^{\prime\prime}}+\widetilde{\Delta}^{\prime\prime}+\lambda_{n}\widetilde{H}^{\prime\prime}).\end{split}$
@@
$g^{*}(K_{X_{n}}+\Delta_{n})=\widetilde{g}^{*}(K_{\widetilde{X}^{\prime\prime}}+\widetilde{\Delta}^{\prime\prime}).$
@@
${(\widetilde{X},\widetilde{\Delta})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${(\widetilde{X}^{\prime},\widetilde{\Delta}^{\prime})}$${(X,\Delta)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(X_{n},\Delta_{n})}$
@@
By Step 4, after shrinking $Y$ around $W$, there is a weak lc model of $(\widetilde{X}^{\prime},\widetilde{\Delta}^{\prime}+\lambda^{\prime}\widetilde{H}^{\prime})$ over $Y$ around $W$. We run a $(K_{\widetilde{X}^{\prime}}+\widetilde{\Delta}^{\prime}+\lambda^{\prime}\widetilde{H}^{\prime})$-MMP over $Y$ around $W$ with scaling of an ample divisor. By Lemma 3.13 and the conclusion of Step 5, the MMP terminates with a log minimal model $(\widetilde{X}^{\prime\prime\prime},\widetilde{\Delta}^{\prime\prime\prime}+\lambda^{\prime}\widetilde{H}^{\prime\prime\prime})$ over $Y$ around $W$. By (i) in Step 3 and shrinking $Y$ around $W$, we get a bimeromorphic contraction

### v2

@@ 3.2. Definition of log MMP
    • $\phi\colon X\dashrightarrow X^{\prime}$ is a bimeromorphic contraction over $Y$ and $f\colon X\to Z$ and $f^{\prime}\colon X^{\prime}\to Z$ are bimeromorphisms,
    • $f$ is a contraction of a $(K_{X}+\Delta)$-negative extremal ray of $\overline{\rm NE}(X/Y;W)$, in particular, $\rho(X/Y;W)-\rho(Z/Y;W)=1$ (cf. [F22b, Theorem 1.2 (4)(iii)]) and $(K_{X}+\Delta)$ is ample over $Z$, and
@@
    For the simplicity of notation, a sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ is denoted by (X_0,Δ_0)⇢(X_1,Δ_1)⇢⋯⇢(X_i,Δ_i)⇢⋯.
@@
    Let $H$ be an $\mathbb{R}$-Cartier divisor on $X$. We put H_0:=H—_X_0   and   H_i+1:=(ϕ_i*H_i)—_X_i+1 for each $i\geq 0$. Then a sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ with scaling of $H$ is a sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ (X_0,Δ_0)⇢(X_1,Δ_1)⇢⋯⇢(X_i,Δ_i)⇢⋯ with the data $\{Y_{i}\}_{i\geq 0}$ and $\{\phi_{i}\colon X_{i}\dashrightarrow X^{\prime}_{i}\}_{i\geq 0}$ such that $(X_{i},\Delta_{i})\dashrightarrow(X^{\prime}_{i},\phi_{i*}\Delta_{i})$ is a step of a $(K_{X_{i}}+\Delta_{i})$-MMP over $Y$ around $W$ with scaling of $H_{i}$ for every $i$.
@@
be a sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ defined with $\{Y_{i}\}_{i\geq 0}$ and $\{\phi_{i}\colon X_{i}\dashrightarrow X^{\prime}_{i}\}_{i\geq 0}$. We say that the $(K_{X}+\Delta)$-MMP is represented by bimeromorphic contractions over $Y$ if $Y_{i}=Y$ for all $i\geq 0$.
@@ 3.3. Basics of log MMP
${(X,\Delta)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$$\scriptstyle{f}$${(X^{\prime},\Delta^{\prime}:=\phi_{*}\Delta)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}}$${Z}$
@@
By the ampleness of $(K_{X}+\Delta)$ over $Z$ and [F22b, Theorem 9.1], after shrinking $Y$ around $W$, we can find $m_{0}\in\mathbb{Z}_{>0}$ such that $mD$ is base point free for all $m\geq m_{0}$. Let $X\to Z_{1}$ and $X\to Z_{2}$ be the contractions over $Z$ induced by $(m_{0}+1)D$ and $(m_{0}+2)D$, respectively. Since $D\equiv_{W}0$, the induced morphisms $Z_{1}\to Z$ and $Z_{2}\to Z$ are biholomorphisms near the inverse image of $W$. By shrinking $Y$ around $W$ if necessary, we may assume $Z_{1}=Z_{2}=Z$. Then there are Cartier divisors $H_{1}$ and $H_{2}$ on $Z$ such that $(m_{0}+1)D\sim f^{*}H_{1}$ and $(m_{0}+2)D\sim f^{*}H_{2}$. Then
@@
${(X,\Delta)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$$\scriptstyle{f}$${(X^{\prime},\Delta^{\prime}:=\phi_{*}\Delta)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}}$${Z}$
@@
Since $\rho(X/Y;W)-\rho(Z/Y;W)=1$ and $(K_{X}+\Delta)$ is ample over $Z$, by shrinking $Y$ around $W$, we can find $r\in\mathbb{R}$ and an $\mathbb{R}$-Cartier divisor $L$ on $Z$ such that
@@ 3.4. Nef thresholds in log MMP
${(X_{i},\Delta_{i})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{i}}$$\scriptstyle{f_{i}}$${(X^{\prime}_{i},\phi_{i*}\Delta_{i})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}_{i}}$${Z_{i}}$
@@
such that $D_{j}$ are Cartier and trivially intersects any curve over $W$ contracted by $f_{i}$. By the ampleness of $(K_{X_{i}}+\Delta_{i})$ over $Z_{i}$ and [F22b, Theorem 9.1], after shrinking $Y$ around $W$, it follows that $D_{j}$ are semi-ample over $Z_{i}$ for all $j$. Let $X_{i}\to V_{j}$ be the contractions over $Z_{i}$ induced by $D_{j}$. By the above condition for $D_{j}$, the induced morphisms $V_{j}\to Z_{i}$ are biholomorphisms near the inverse image of $W$. By shrinking $Y$ around $W$, we may assume $V_{j}=Z_{i}$ for all $j$. Then there are $\mathbb{Q}$-Cartier divisors $H_{j}$ on $Z_{i}$ such that $D_{j}\sim_{\mathbb{Q}}f_{i}^{*}H_{j}$ for all $j$. Put $A:=\sum_{j}r_{j}H_{j}$. Then
@@ 3.5. From existence of weak lc model to termination
(*) $\begin{split}K_{\widetilde{X}}+\widetilde{\Delta}+t\widetilde{H}=f^{*}(K_{X}+\Delta+tH)+\widetilde{E}_{t}.\end{split}$
@@
[…]
@@
[…]
@@
$1=a(P,X^{\prime},\Delta^{\prime})\geq a(P,X,\Delta)\geq-1,$
@@
[…]
@@
[…]
@@
By Lemma 3.14, the induced bimeromorphic map $\widetilde{f}^{\prime}\colon\widetilde{X}^{\prime}\dashrightarrow X^{\prime}$ is a morphism. Let $\widetilde{E}^{\prime}$ and $\widetilde{F}^{\prime}$ be the strict transforms of $\widetilde{E}_{0}$ and $\widetilde{F}$ on $\widetilde{X}^{\prime}$, respectively. Then we have $\widetilde{L}^{\prime}\sim_{\mathbb{R}}(3\cdot{\rm dim}\,X)\widetilde{f}^{\prime*}L^{\prime}$, and (**1) implies
@@
[…]
@@
[…]
@@
[…]
@@
is represented by bimeromorphic contractions over $Y$ and $X=X_{0}$.
@@
[…]
@@
    $f\colon\widetilde{X}\to X$ is a log resolution of $(X,\Delta+H)$ such that for any $t\in[0,1]$, we can write K_~X+~Δ+t~H=f^*(K_X+Δ+tH)+~E_t for some effective $f$-exceptional $\mathbb{R}$-divisor $\widetilde{E}_{t}$ on $\widetilde{X}$, where $\widetilde{H}=f^{-1}_{*}H$,
@@
    $\widetilde{X}\dashrightarrow\widetilde{X}^{\prime}$ is a bimeromorphic contraction over $Y$ that defines:
[…]
@@
[…]
@@
    $X\dashrightarrow X_{n}$ is a bimeromorphic contraction over $Y$ that defines a finite sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ with scaling of $H$ (X, Δ) ⇢(X_n, Δ_n) such that $\lambda_{n}<{\rm min}\{\lambda_{n-1},\epsilon\}$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]